# 3DRot: 3D Rotation Augmentation for RGB-Based 3D Tasks

### v1

@@ Introduction
Random horizontal flipping is the most widely used geometry-aware transform in RGB-based tasks because it is label‐preserving for the image plane and inexpensive to implement. The pioneering (Brazil and Liu 2019) pipeline reflects each 3D box corner on the image, then back-projects the flipped vertices with the original depth to obtain an updated 3D cuboid, leaving all other pose parameters. Although effective for heading predictions, this strategy cannot propagate sign changes to the off-diagonal entries of a full ${\times}3$ 3 rotation matrix and therefore assumes that only the yaw angle is modeled. Several follow-up methods (Chen et al. 2020; Zou et al. 2021; Li et al. 2022; Li, Jia, and Shi 2024; Wang et al. 2021; Pan et al. 2025) inherit the same augmentation. Cube R-CNN (Brazil et al. 2023) introduces a more principled update: it composes two mirror matrices, one for the image warp and one for the camera coordinate frame, so that the predicted $R\!\in\!SO(3)$ remains valid after a flip. This design finally enables horizontal mirroring for full‐angle pose regression, but the derivation presumes that the principal point coincides with the image center; when intrinsics are asymmetric or cropped, the 2D-3D alignment still drifts.
@@
denote the 3×83$\times$ 83 × 8 matrix of the eight cuboid corners in the original camera coordinates, where the column matrix $S_{bb}\in R^{3\times 8}$ is formed by concatenating the vectorized offsets (±$\bigl{(}\pm\frac{W}{2},\,\pm\frac{H}{2},\,\pm\frac{L}{2}\bigr{)}^{\!\top}$.
@@
It is important to note that $P_{A}$ and $P_{B}$, while elements of $\mathbf{R}^{3\times 1}$, represent points on their respective image planes in homogeneous coordinates, typically of the form $[u,v,1]^{T}$. Given that the third component of these homogeneous image coordinates is inherently 1 (after normalization), the scaling factor $\frac{z_{B}}{z_{A}}$ in Eq. (8) can be conceptualized as a normalization constant denoted by $\lambda$. This scalar $\lambda$ ensures that the resulting vector $P_{A}$ is correctly scaled to its homogeneous representation where the last component is 1. Thus, Eq. (8) can be expressed in a more general form, highlighting the projective transformation up to a scale factor:
@@
Setting $O_{1}=I$ yields $A^{\!\top}A=D^{2}$. Now recall the uniqueness theorem of second-order isotropic tensors (Reddy 2018) : If a real symmetric tensor T satisfies $Q\,T\,Q^{\top}=T$ for all $Q\in O(3)$, then $T=\lambda I$
[…]
@@
Screen coordinates place the origin at the top-left corner (0,0)(0,0)( 0 , 0 ), whereas camera intrinsics locate the principal point $(c_{x},c_{y})$ near the image center. After rotation, points that were inside the frame can land at negative coordinates if projected with the original intrinsics. Translating the image to make all pixels positive and then cropping back to the original size therefore shifts the geometric center away from the declared principal point, corrupting downstream geometry.
@@
    Project the original corners. Let $P_{\mathrm{orig}}$ be the four image corners and $R$ the applied camera rotation. Define $K_{0}$ to equal the source intrinsics $K_{A}$ but with principal point (0,0)(0,0)( 0 , 0 ). The corners in the rotated frame are
@@
    Update intrinsics and crop. Build a new intrinsic matrix $K_{C}$ equal to the source intrinsics $K_{A}$ but with principal point $(x_{\max},y_{\max})$. Render the rotated image with $K_{C}$ and finally crop the rectangle $[0,2x_{\max}]\times[0,2y_{\max}]$.
@@
[…]
@@
Table 1: Single-run results on SUN10 and IN10. All models share the same frozen DINO-X + Cube R-CNN backbone and training schedule. 3DRot applies a camera-centric rotation (with probability 0.8; yaw ±$\pm 10^{\circ}$, pitch/roll ±$\pm 5^{\circ}$) plus a chirality-preserving horizontal flip with probability 0.5; rotation and flip are composable.
@@
The last decoder layer produces $N{=}900$ object queries (256 dims). We feed the corresponding image features into a Cube R-CNN head (Brazil et al. 2023), which predicts a 9-D cuboid representation-2D center, depth, 3D dimensions, and a full ${\times}3$ 3 rotation matrix.
@@
Table 1 summarizes single-run performance (seed 42) on the two evaluation splits. The first two rows correspond to the SUN10 setting: a vanilla DINO-X + Cube R-CNN baseline (labeled SUN10) and the same backbone equipped with 3DRot. Unless stated otherwise, 3DRot is applied with probability 0.8 for an in-plane rotation (yaw ±$\pm 10^{\circ}$, pitch/roll ±$\pm 5^{\circ}$) and probability 0.5 for a chirality-preserving horizontal flip; flips may be composed with the rotation within the same image.
@@
Padding the rotated image and realigning the principal point (column Center) brings a small but consistent gain: removing it degrades $IoU_{3D}$ from 44.58 to 44.43 and leaves $mAP_{0.5}$ unchanged. Chirality preservation is critical: turning it off inflates the rotation error from 20.78° to 23.57° and lowers $mAP_{0.5}$ by 0.57, confirming that naively mirroring a full $\times 3$ 3 rotation matrix breaks orientation consistency.

### v2

@@ Abstract
11footnotetext: This work was done during an internship at the International Digital Economy Academy (IDEA).33footnotetext: Corresponding author.
@@ Introduction
Random horizontal flipping is the most widely used geometry-aware transform in RGB-based tasks because it is label‐preserving for the image plane and inexpensive to implement. The pioneering (Brazil and Liu 2019) pipeline reflects each 3D box corner on the image, then back-projects the flipped vertices with the original depth to obtain an updated 3D cuboid, leaving all other pose parameters. Although effective for heading predictions, this strategy cannot propagate sign changes to the off-diagonal entries of a full $3{\times}3$ rotation matrix and therefore assumes that only the yaw angle is modeled. Several follow-up methods (Chen et al. 2020; Zou et al. 2021; Li et al. 2022; Li, Jia, and Shi 2024; Wang et al. 2021; Pan et al. 2025) inherit the same augmentation. Cube R-CNN (Brazil et al. 2023) introduces a more principled update: it composes two mirror matrices, one for the image warp and one for the camera coordinate frame, so that the predicted $R\!\in\!SO(3)$ remains valid after a flip. This design finally enables horizontal mirroring for full‐angle pose regression, but the derivation presumes that the principal point coincides with the image center; when intrinsics are asymmetric or cropped, the 2D-3D alignment still drifts.
@@
denote the $3\times 8$ matrix of the eight cuboid corners in the original camera coordinates, where the column matrix $S_{bb}\in R^{3\times 8}$ is formed by concatenating the vectorized offsets $\bigl{(}\pm\frac{W}{2},\,\pm\frac{H}{2},\,\pm\frac{L}{2}\bigr{)}^{\!\top}$.
@@
It is important to note that $P_{A}$ and $P_{B}$, while elements of $\mathbf{R}^{3\times 1}$, represent points on their respective image planes in homogeneous coordinates, typically of the form $u,v,1]^{T}$. Given that the third component of these homogeneous image coordinates is inherently 1 (after normalization), the scaling factor $\frac{z_{B}}{z_{A}}$ in Eq. (8) can be conceptualized as a normalization constant denoted by $\lambda$. This scalar $\lambda$ ensures that the resulting vector $P_{A}$ is correctly scaled to its homogeneous representation where the last component is 1. Thus, Eq. (8) can be expressed in a more general form, highlighting the projective transformation up to a scale factor:
@@
Setting $O_{1}=I$ yields $A^{\!\top}A=D^{2}$. Now recall the uniqueness theorem of second-order isotropic tensors (Reddy 2018) : If a real symmetric tensor T satisfies $Q\,T\,Q^{\top}=T$ for all $Q\in O(3)$, then $T=\lambda I$ .
Applying this to the symmetric tensor $T=A^{\!\top}A$ gives $A^{\!\top}A=\lambda^{2}I$, hence $A=\lambda R_{A}$ for some $\lambda>0$ and $R_{A}\in\mathrm{O}(3)$. In the context of axis-aligned cuboids we further fix $R_{A}=I$, so the only admissible affine map is a uniform scaling $A=\lambda I$. Affine transforms with shears or non-uniform scalings therefore cannot preserve the $(R,t,s)$ factorization for all poses and are left to the Discussion section for domain-specific treatment.
@@
Screen coordinates place the origin at the top-left corner $(0,0)$, whereas camera intrinsics locate the principal point $(c_{x},c_{y})$ near the image center. After rotation, points that were inside the frame can land at negative coordinates if projected with the original intrinsics. Translating the image to make all pixels positive and then cropping back to the original size therefore shifts the geometric center away from the declared principal point, corrupting downstream geometry.
@@
    Project the original corners. Let $P_{\mathrm{orig}}$ be the four image corners and $R$ the applied camera rotation. Define $K_{0}$ to equal the source intrinsics $K_{A}$ but with principal point $(0,0)$. The corners in the rotated frame are
@@
    Update intrinsics and crop. Build a new intrinsic matrix $K_{C}$ equal to the source intrinsics $K_{A}$ but with principal point $(x_{\max},y_{\max})$. Render the rotated image with $K_{C}$ and finally crop the rectangle $0,2x_{\max}]\times[0,2y_{\max}]$.
@@
We report seven metrics taken from Cube R-CNN and NOCSFormer (Brazil et al. 2023; Krishnan et al. 2024): $IoU_{3D}$, TRANS [cm], ROT [deg], SIZE [cm], 5GRAV, 10HEAD and $mAP_{0.5}$. $IoU_{3D}$ is computed with the exact volumetric intersection between predicted and ground-truth cuboids, not the fast approximation used in the original Cube R-CNN code. AP denotes mean Average Precision at a fixed 2D IoU threshold of 0.5. $mAP_{0.5}$ averages AP over the nine 3D IoU thresholds $\{0.05,0.10,\dots,0.50\}$, providing a finer-grained view of performance. 5GRAV is the percentage of predictions with gravity-axis error $5^{\circ}$; 10HEAD is the percentage with heading (yaw) error $10^{\circ}$ (Krishnan et al. 2024). For $IoU_{3D}$, 5GRAV, 10HEAD and $mAP_{0.5}$ higher values indicate better performance, whereas lower values are preferred for TRANS, ROT, and SIZE.
@@
Table 1: Single-run results on SUN10 and IN10. All models share the same frozen DINO-X + Cube R-CNN backbone and training schedule. 3DRot applies a camera-centric rotation (with probability 0.8; yaw $\pm 10^{\circ}$, pitch/roll $\pm 5^{\circ}$) plus a chirality-preserving horizontal flip with probability 0.5; rotation and flip are composable.
@@
The last decoder layer produces N=900N{=}900 object queries (256 dims). We feed the corresponding image features into a Cube R-CNN head (Brazil et al. 2023), which predicts a 9-D cuboid representation-2D center, depth, 3D dimensions, and a full $3{\times}3$ rotation matrix.
@@
Table 1 summarizes single-run performance (seed 42) on the two evaluation splits. The first two rows correspond to the SUN10 setting: a vanilla DINO-X + Cube R-CNN baseline (labeled SUN10) and the same backbone equipped with 3DRot. Unless stated otherwise, 3DRot is applied with probability 0.8 for an in-plane rotation (yaw $\pm 10^{\circ}$, pitch/roll $\pm 5^{\circ}$) and probability 0.5 for a chirality-preserving horizontal flip; flips may be composed with the rotation within the same image.
@@
Padding the rotated image and realigning the principal point (column Center) brings a small but consistent gain: removing it degrades $IoU_{3D}$ from 44.58 to 44.43 and leaves $mAP_{0.5}$ unchanged. Chirality preservation is critical: turning it off inflates the rotation error from 20.78° to 23.57° and lowers $mAP_{0.5}$ by 0.57, confirming that naively mirroring a full $3\times 3$ rotation matrix breaks orientation consistency.